# Entropic balance with feedback control: information equalities and tight inequalities

### v1

@@ Abstract
Discussion– The improvement on the tightness of second-law inequalities stemming from the Markovian description is appealing from a conceptual point of view: the Markovian description uses the least information needed to analyse the dynamics of the system by only retaining the value of the control variable that determines the potential felt by the system. Also, the Markovian second-law inequalities, as given by Eq. (7), quite accurately predict the range of parameters in which model systems are able to extract work and thus are useful as information engines—especially when the housekeeping contribution to the entropy production is kept. This is a task at which previous information inequalities in the literature, stemming from the chain description, fail.
@@
Ashida et al. introduced the concept of unavailable information to improve the existing bounds for the extractable work [13]. Within our Markovian modelling, we have shown analytically that the unavailable information’s bound saturates for error-free measurements in a wide class of systems; Szilard’s engine [13] and the “lossless” information engine of Refs. [17, 49] are just particular cases, for $\Delta t_{\text{m}}\to\infty$, of the general result (11) derived in this Letter. Furthermore, we have investigated how the unavailable information behaves under imperfect feedback protocols, a setting that remained largely unexplored. Interestingly, the unavailable information bound is no longer saturated and, in fact, there exist regions of measurement error in which the Markovian bound, $\Delta_{\text{m}}I$, improves the unavailable information’s, $\overline{\mathcal{I}_{u}-\mathcal{I}_{\vec{x}}}$.
@@
Information functions appear in the second law as a consequence of the implicit coarse-graining carried out when discarding the internal state of the feedback controller in the description of the system plus controller. Introducing a microscopic modelling of the internal state of the feedback controller and its interaction with the system would make it possible to identify the microscopic origin of the information function in the mesoscopic second law. This is a relevant question that deserves to be investigated. On the other hand, the framework developed here applies to overdamped systems, a typical experimental situation for information engines. Extending it to the underdamped regime [50, 51, 52], which takes into account inertial effects, is another appealing perspective.
@@
Data Availability Statement– The codes employed for generating the data are openly available in the GitHub page of University of Sevilla’s FINE research group.

### v2

@@ Abstract
Discussion– The improvement on the tightness of second-law inequalities stemming from the Markovian description is appealing from a conceptual point of view: the Markovian description uses the least information needed to analyse the system’s dynamics by only retaining the value of the control variable determining the potential felt by the system. Also, the Markovian second-law inequalities, as given by Eq. (8), quite accurately predict the range of parameters in which model systems are able to extract work and thus are useful as information engines—for non-equilibium situations in which detailed balance is broken, the housekeeping contribution to the entropy production should be retained. This is a task at which previous information inequalities in the literature, stemming from the chain description, fail.
@@
Ashida et al. introduced the concept of unavailable information $\overline{\mathcal{I}_{u}}$ to improve the existing bounds for the extractable work [13]. Within our Markovian modelling, we have shown analytically that the unavailable information’s bound saturates for error-free measurements in a wide class of systems; Szilard’s engine [13] and the “lossless” information engine of Refs. [15, 52]—see also End Matter—are just particular cases, for $\Delta t_{\text{m}}\to\infty$, of the general result (12) derived in this Letter. Furthermore, we have investigated how $\overline{\mathcal{I}_{u}}$ behaves under imperfect feedback protocols, a setting that remained largely unexplored. Interestingly, the unavailable information bound is no longer saturated and, in fact, there exist regions of measurement error in which the Markovian bound, $\Delta_{\text{m}}I$, improves the unavailable information’s, $\overline{\mathcal{I}_{u}-\mathcal{I}_{\vec{x}}}$.
The Markovian framework accurately describes systems with feedback control in many experimental situations, when the controller’s time delay is negligible against the system’s typical relaxation time in contact with the thermal bath. Otherwise, the Markovian bound would be no longer applicable and one has to resort to either the transfer entropy bound or the unavailable information bound to estimate the extractable work.
A key consideration is the choice of variables on which the information bounds depend. We have used the system state $x$ and the control variable $c$, which is updated at each measurement with probability $\Theta(c_{n}|x_{n})$, to obtain the bound $\Delta_{\text{m}}I(x,c)$. An alternative approach uses $x$, a measurement outcome $x^{\prime}$—e.g. $x$ plus measurement noise—and a deterministic control update $c_{n}=f(x^{\prime}_{n})$, generating the same dynamics. When only the noisy outcome $x^{\prime}$ is available, the Markovian bound $\Delta_{\text{m}}I(x^{\prime},c)$ still holds, but this bound systematically deteriorates as $x^{\prime}$ becomes more coarse-grained.
@@
Information functions in the second law stem from the implicit coarse-graining carried out when discarding the internal state of the feedback controller in the description of the system plus controller. Introducing a microscopic modelling of the internal state of the feedback controller and its interaction with the system would make it possible to identify the microscopic origin of the information function in the mesoscopic second law. This is a relevant question that deserves to be investigated. On another note, the framework developed here applies to overdamped information engines. Extending it to the underdamped regime [53, 54, 55], which takes into account inertial effects, is another appealing perspective.
@@
Data Availability Statement– The codes employed for generating the data are openly available on the GitHub page of University of Seville’s FINE research group [56].